A market research firm conducts telephone surveys with a 28% historical response rate. What is the probability that in a new sample of 600 telephone numbers, at least 150 individuals will cooperate and respond to the questions? In other words, what is the probability that the sample proportion will be at least

150/600  = 0.25?
Answer #1

n = Sample Size = 600

p = Population Proportion = 0.28

q = 1 - p = 0.72

SE =

To find P( 0.25):

Z= (0.25 - 0.28)/0.0183

= - 1.6366

By Technology, Cumulative Area Under Standard Normal Curve = 0.0509

So,

( 0.25):= 1 - 0.0509 = 0.9491

So,

Answer is:

0.9491